Given that 4x/5 – 1/10 ≤ x –2/5, state the smallest value of x when:
i) x is an integer
ii) x is a prime number
iii) x is a rational number

Answer by RAY100(1637)

(4x)/5 -1/10 <= x-2/5
.
multiply by 10
.
8x -1 <= 10x -4
.
3 <= 2x
.
3/2 <=x,,,,,or x>= 1.5
.
i),,,, 2,,smallest integer > 1.5
.
ii),,, 2,, smallest prime > 1.5
.
iii),,,3/2, smallest rational number >=1.5
.